b1az3:

Match each function with the point that represents the vertex of its graph. No one point be used more than once.

b1az3:

1 attachment
darkknight:

so if a function is written in this form: \[y=ax^2+bx+c\] the x coordinate of the vertex is \[-b/2a\] and you can take this x coordinate and plug back into the equation to get the y coordinate

darkknight:

that is if a function is written in standard form ^

darkknight:

if a function is written in vertex form \[y=a(x-h)^2+k\] (h, k) is the vertex

darkknight:

for your number 2: \[y=-2(x+2)^2\] there is no k so k = 0 h = -2 here do you get this?

darkknight:

for your number 5:\[y=-2x^2+4x+2\] find -b/2a, where in this case a= -2 and b = 4, thats the x coordinate of the vertex then plug in the x coordinate into the function to find the y-coordinate

darkknight:

If you have something in this form \[(x+1)(x+2)\] you can use FOIL and factor this out into x^2 + 2x + x + 2 = x^2 +3x+2

darkknight:

@b1az3 still here?

b1az3:

yes

offbeatsoldier333334:

how do you make your car on gta drive under water

b1az3:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @offbeatsoldier333334 how do you make your car on gta drive under water \(\color{#0cbb34}{\text{End of Quote}}\) wrong site or post

offbeatsoldier333334:

ok